OpenStudy (anonymous):

is the sequence geometric? if so, identify the common. 2, 6, 18, 54,...

OpenStudy (mertsj):

You can tell if it is geometric by dividing each term by the previous term.

OpenStudy (mertsj):

For example: 6/2=3 18/6=3 54/18=3

OpenStudy (mertsj):

So yes...it is geometric because we always get the same thing. And the number we get, the 3, is called the common ratio.

OpenStudy (anonymous):

@Mertsj so, the identity common is?

OpenStudy (mertsj):

The common ratio which is 3, r = 3

OpenStudy (mertsj):

Do you mean: \[a _{n}=ra _{n-1}\]

OpenStudy (anonymous):

you are perfect